OpenStudy (anonymous):

Which of the following is an equation of the line that passes through the point (1, 1) and has a slope of 2? A. y = 2x – 1 B. y = 2x + 1 C. y = –2x – 1 D. y = –2x + 1

OpenStudy (zzr0ck3r):

we want the form y = mx+b m = slope = 2 y = 2x+b we need to find b and we are done use the point (1,1) so x = 1, y = 1 y = 2x+ b becomes 1 = 2(1) + b b = -1 so what is the answer?

OpenStudy (zale101):

Start with the slope formula.

OpenStudy (anonymous):

B

OpenStudy (zzr0ck3r):

y = 2x+ b b = -1 what is the answer?

OpenStudy (anonymous):

C

OpenStudy (zzr0ck3r):

put -1 where you see b in the following equation y=2x+b and then type me what you have

OpenStudy (anonymous):

TO TIRED TO HINK ....

OpenStudy (anonymous):

THATS not a option

OpenStudy (zzr0ck3r):

y=2x+b ^put -1 there

OpenStudy (anonymous):

y=2x+-1

OpenStudy (zzr0ck3r):

ok what is x+(-1)?

OpenStudy (zzr0ck3r):

2x-1 is the answer

OpenStudy (zzr0ck3r):

gl

OpenStudy (anonymous):

Actually we dont need any fromula here just plug the x any y in equation and voila!